An open box is to be made from a square sheet of tin-plate by cutting out 2 cm squares from each corner and folding up the sides. If the required capacity of the box is to be 32cm3, find the dimensions of the sheet of the tin-plate required.
Answer by mananth(16946)

let the side of the sheet be x
2 cm cut off from both ends
side of box = (x-4)
height = 2cm.
Volume of box = base ^2 * height
32= (x-4)^2*2
16= (x-4)^2
take square root
x-4 = 4
x= 8
Side of sheet = 8 cm